OpenStudy (erinweeks):

HELPP Suppose that a sphere with radius 5a has the same volume as a cone of radius 3a. What is the height of the cone? Give your answer in terms of a. Use the formulas and for the volumes of a sphere and cone respectively.

OpenStudy (erinweeks):

FORMULAS : \[V=\frac{ 4 }{ 3 } \pi r ^{3}\] \[V = \frac{ 1 }{ 3 }\pi r ^{2} h\]

hartnn (hartnn):

just put r=5a in 1st equation r=3a in 2nd equation. then divide 1st equation by 2nd equation. u will get h after simplification.

OpenStudy (erinweeks):

i got \[\frac{ 20 }{ 3 }\]

hartnn (hartnn):

nopes.....also there should be 'a' in your answer

OpenStudy (erinweeks):

answers .. b. \[\frac{ 500 }{ 9 }a\] c. \[\frac{ 9 }{ 500 }a ^{2}\] d. \[\frac{ 500 }{ 9 }a\]

OpenStudy (erinweeks):

these are the possible answers

hartnn (hartnn):

\[4(125a^3)=(9a^2)h\] so \[h=\frac{125*4*a^3}{9a^2}\]

hartnn (hartnn):

can u solve further?

hartnn (hartnn):

whats the difference between option b and d ??

OpenStudy (erinweeks):

b is suppose to be a squared sorry !

hartnn (hartnn):

so can u get h from \[h=\frac{125*4*a^3}{9a^2}\] ?

OpenStudy (erinweeks):

wouldnt it be d ? i think ...

hartnn (hartnn):

125*4=500

hartnn (hartnn):

yup :)

OpenStudy (erinweeks):

so its d i got it right

hartnn (hartnn):

yes,its d

OpenStudy (erinweeks):

thank you so much for the help !